Measuring Program Resource Use

The GNU time Command Edition 1.7, for version 1.7

12 June 1996

David MacKenzie

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1991, 92, 93, 96 Free Software Foundation, Inc.

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies.

Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one.

Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the Foundation.

1 Measuring Program Resource Use

The time command runs another program, then displays information about the resources used by that program, collected by the system while the program was running. You can select which information is reported and the format in which it is shown (see Section 1.1 [Setting Format], page 1), or have time save the information in a file instead of displaying it on the screen (see Section 1.3 [Redirecting], page 4).

The resources that time can report on fall into the general categories of time, memory, and I/O and IPC calls. Some systems do not provide much information about program resource use; time reports unavailable information as zero values (see Section 1.5 [Accuracy], page 6).

The format of the time command is:

time [option…] command [arg…]

time runs the program command, with any given arguments arg. . .. When command finishes, time displays information about resources used by command.

Here is an example of using time to measure the time and other resources used by running the program grep:

     eg$ time grep nobody /etc/aliases
     nobody:/dev/null
     etc-files:nobody
     misc-group:nobody
     0.07user 0.50system 0:06.69elapsed 8%CPU (0avgtext+489avgdata 324maxresident)k
     46inputs+7outputs (43major+251minor)pagefaults 0swaps

Mail suggestions and bug reports for GNU time to bug-gnu-utils@prep.ai.mit.edu. Please include the version of time, which you can get by running ‘time --version’, and the operating system and C compiler you used.

1.1 Setting the Output Format

time uses a format string to determine which information to display about the resources used by the command it runs. See Section 1.2 [Format String], page 2, for the interpretation of the format string contents.

You can specify a format string with the command line options listed below. If no format is specified on the command line, but the TIME environment variable is set, its value is used as the format string. Otherwise, the default format built into time is used:

     %Uuser %Ssystem %Eelapsed %PCPU (%Xtext+%Ddata %Mmax)k
     %Iinputs+%Ooutputs (%Fmajor+%Rminor)pagefaults %Wswaps

The command line options to set the format are:

-f format

--format=format

Use format as the format string.

-p
--portability

Use the following format string, for conformance with POSIX standard 1003.2:

               real %e
               user %U
               sys %S
-v
--verbose

Use the built-in verbose format, which displays each available piece of information on the program’s resource use on its own line, with an English description of its meaning.

1.2 The Format String

The format string controls the contents of the time output. It consists of resource specifiers and escapes, interspersed with plain text.

A backslash introduces an escape, which is translated into a single printing character upon output. The valid escapes are listed below. An invalid escape is output as a question mark followed by a backslash.

\t a tab character

\n a newline

\\ a literal backslash

time always prints a newline after printing the resource use information, so normally format strings do not end with a newline character (or ‘\n’).

A resource specifier consists of a percent sign followed by another character. An invalid resource specifier is output as a question mark followed by the invalid character. Use ‘%%’ to output a literal percent sign.

The resource specifiers, which are a superset of those recognized by the tcsh builtin time command, are listed below. Not all resources are measured by all versions of Unix, so some of the values might be reported as zero (see Section 1.5 [Accuracy], page 6).

1.2.1 Time Resources

E Elapsed real (wall clock) time used by the process, in [hours:]minutes:seconds. e Elapsed real (wall clock) time used by the process, in seconds. S Total number of CPU-seconds used by the system on behalf of the process (in kernel

mode), in seconds.

U Total number of CPU-seconds that the process used directly (in user mode), in seconds. P Percentage of the CPU that this job got. This is just user + system times divied by

the total running time.

1.2.2 Memory Resources

M Maximum resident set size of the process during its lifetime, in Kilobytes. t Average resident set size of the process, in Kilobytes. K Average total (data+stack+text) memory use of the process, in Kilobytes. D Average size of the process’s unshared data area, in Kilobytes. p Average size of the process’s unshared stack, in Kilobytes. X Average size of the process’s shared text, in Kilobytes. Z System’s page size, in bytes. This is a per-system constant, but varies between systems.

1.2.3 I/O Resources

F Number of major, or I/O-requiring, page faults that occurred while the process was running. These are faults where the page has actually migrated out of primary memory.

R Number of minor, or recoverable, page faults. These are pages that are not valid (so they fault) but which have not yet been claimed by other virtual pages. Thus the data in the page is still valid but the system tables must be updated.

W Number of times the process was swapped out of main memory.

c Number of times the process was context-switched involuntarily (because the time slice expired).

w Number of times that the program was context-switched voluntarily, for instance while waiting for an I/O operation to complete.

I Number of file system inputs by the process.

O Number of file system outputs by the process.

r Number of socket messages received by the process.

s Number of socket messages sent by the process.

k Number of signals delivered to the process.

1.2.4 Command Info

C Name and command line arguments of the command being timed.

x Exit status of the command.

1.3 Redirecting Output

By default, time writes the resource use statistics to the standard error stream. The options below make it write the statistics to a file instead. Doing this can be useful if the program you’re running writes to the standard error or you’re running time noninteractively or in the background.

-o file

--output=file

Write the resource use statistics to file. By default, this overwrites the file, destroying the file’s previous contents.

-a

--append Append the resource use information to the output file instead of overwriting it. This option is only useful with the ‘-o’ or ‘--output’ option.

1.4 Examples

Run the command ‘wc /etc/hosts’ and show the default information:

eg$ time wc /etc/hosts
      35     111    1134 /etc/hosts
0.00user 0.01system 0:00.04elapsed 25%CPU (0avgtext+0avgdata 0maxresident)k
1inputs+1outputs (0major+0minor)pagefaults 0swaps

Run the command ‘ls -Fs’ and show just the user, system, and wall-clock time:

eg$ time -f "\t%E real,\t%U user,\t%S sys" ls -Fs
total 16
1 account/      1 db/           1 mail/         1 run/
1 backups/      1 emacs/        1 msgs/         1 rwho/
1 crash/        1 games/        1 preserve/     1 spool/
1 cron/         1 log/          1 quotas/       1 tmp/
        0:00.03 real,   0.00 user,      0.01 sys

Edit the file ‘.bashrc’ and have time append the elapsed time and number of signals to the file ‘log’, reading the format string from the environment variable TIME:

eg$ export TIME="\t%E,\t%k" # If using bash or ksh
eg$ setenv TIME "\t%E,\t%k" # If using csh or tcsh
eg$ time -a -o log emacs .bashrc
eg$ cat log
        0:16.55,        726

Run the command ‘sleep 4’ and show all of the information about it verbosely:

eg$ time -v sleep 4
        Command being timed: "sleep 4"
        User time (seconds): 0.00
        System time (seconds): 0.05
        Percent of CPU this job got: 1%
        Elapsed (wall clock) time (h:mm:ss or m:ss): 0:04.26
        Average shared text size (kbytes): 36

Average unshared data size (kbytes): 24
Average stack size (kbytes): 0
Average total size (kbytes): 60
Maximum resident set size (kbytes): 32
Average resident set size (kbytes): 24
Major (requiring I/O) page faults: 3
Minor (reclaiming a frame) page faults: 0
Voluntary context switches: 11
Involuntary context switches: 0
Swaps: 0
File system inputs: 3
File system outputs: 1
Socket messages sent: 0
Socket messages received: 0
Signals delivered: 1
Page size (bytes): 4096
Exit status: 0

1.5 Accuracy

The elapsed time is not collected atomically with the execution of the program; as a result, in bizarre circumstances (if the time command gets stopped or swapped out in between when the program being timed exits and when time calculates how long it took to run), it could be much larger than the actual execution time.

When the running time of a command is very nearly zero, some values (e.g., the percentage of CPU used) may be reported as either zero (which is wrong) or a question mark.

Most information shown by time is derived from the wait3 system call. The numbers are only as good as those returned by wait3. Many systems do not measure all of the resources that time can report on; those resources are reported as zero. The systems that measure most or all of the resources are based on 4.2 or 4.3BSD. Later BSD releases use different memory management code that measures fewer resources.

On systems that do not have a wait3 call that returns status information, the times system call is used instead. It provides much less information than wait3, so on those systems time reports most of the resources as zero.

The ‘%I’ and ‘%O’ values are allegedly only “real” input and output and do not include those supplied by caching devices. The meaning of “real” I/O reported by ‘%I’ and ‘%O’ may be muddled for workstations, especially diskless ones.

1.6 Running the time Command

The format of the time command is:

time [option…] command [arg…]

time runs the program command, with any given arguments arg. . .. When command finishes, time displays information about resources used by command (on the standard error output, by default). If command exits with non-zero status or is terminated by a signal, time displays a warning message and the exit status or signal number.

Options to time must appear on the command line before command. Anything on the command line after command is passed as arguments to command.

-o file

--output=file

Write the resource use statistics to file.

-a

--append Append the resource use information to the output file instead of overwriting it.

-f format

--format=format

Use format as the format string.

--help Print a summary of the command line options to time and exit.

-p
--portability

Use the POSIX format.

-v
--verbose

Use the built-in verbose format.

-V
--version

Print the version number of time and exit.

8 Measuring Program Resource Use

Table of Contents

1 Measuring Program Resource Use . . . . . . . . . . . . . . . 1

1.1 Setting the Output Format. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 The Format String . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

1.2.1 Time Resources . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2.2 Memory Resources. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2.3 I/O Resources . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.2.4 Command Info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

1.3 Redirecting Output. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.4 Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.5 Accuracy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.6 Running the time Command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ii Measuring Program Resource Use